OpenStudy (anonymous):

Medal to the first right answer!!with soln.. Integral of {x^2/(xsinx+cosx)}dx

OpenStudy (anonymous):

\[\int\limits_{?}^{?}{x^2}/(xsinx+cosx)\]

OpenStudy (chosenmatt):

question?

OpenStudy (anonymous):

here you go!!

OpenStudy (chosenmatt):

that is too hard...

OpenStudy (anonymous):

\[\int\limits_{?}^{?}{{x^2}/(xsinx+cosx)}dx\]

OpenStudy (anonymous):

believe me I know

OpenStudy (anonymous):

spent the whole day!!

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

@ParthKohli @dan815 @SolomonZelman

OpenStudy (anonymous):

@Whitemonsterbunny17 @robtobey @.Sam.

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

cmon guys @least give it a try!!

OpenStudy (bohotness):

i got it :)

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

wow..........how??

OpenStudy (anonymous):

Either ur a genius OR ur wrong...

OpenStudy (anonymous):

Shoot!!

OpenStudy (bohotness):

i post it ok

OpenStudy (anonymous):

pls do

OpenStudy (catlover5925):

uhhh im in advanced math and I have never done this...sorry

OpenStudy (bohotness):

eait what grade you in

OpenStudy (anonymous):

eat my grade??........did you solve it??

OpenStudy (anonymous):

I am guessing no!!

OpenStudy (bohotness):

dude or girl i need to known what grade your in or i cant help you

OpenStudy (anonymous):

@kag @Owlfred @LiteNing1337 @Abhisar @skullpatrol @Secret-Ninja @Compassionate @demonchild99 @ganeshie8 @emmigrace222 @mathstudent55

ganeshie8 (ganeshie8):

what makes you think there exists a closed form for this ?

OpenStudy (anonymous):

dont getcha!!

OpenStudy (anonymous):

Is it solvable......I didnt check wolfram!!

OpenStudy (anonymous):

cuz I solved a different but similar one!!

ganeshie8 (ganeshie8):

im not so sure if il able to work an integral that wolfram gave up :O http://www.wolframalpha.com/input/?i=Integral+of+x%5E2%2F%28xsinx%2Bcosx%29+dx

OpenStudy (anonymous):

\[\int\limits_{?}^{?}{ x^2/{(xsinx+cosx)^2}}\]

OpenStudy (anonymous):

lol dude......anyways just wanted to make sure no one else can solve this!!

ganeshie8 (ganeshie8):

@Princer_Jones @perl

OpenStudy (perl):

one sec

OpenStudy (perl):

one sec

OpenStudy (anonymous):

BTW The second one I posted is xtremely tricky....anyone wants the sol??

OpenStudy (perl):

integral x^2 / ( xsin x + cosx) is not an elementary antiderivative

OpenStudy (anonymous):

\[{xsecx.xcosx/{(xsinx+cosx)^2}}\]

OpenStudy (perl):

you want the indefinite integral or the definite integral, i notice you have question marks in your limits

OpenStudy (anonymous):

u have to do like this for the original question

OpenStudy (anonymous):

indefinite

OpenStudy (perl):

this was the original question Integral of {x^2/(xsinx+cosx)}dx ?

OpenStudy (perl):

ok you did sec x * cos x = 1 , i follow so far

OpenStudy (anonymous):

yes...........the one I posted some comments bfore is another form of it which I solved

OpenStudy (perl):

why is there suddenly a square term in the denominator?

OpenStudy (perl):

in your original question there was no square term in the denominator

OpenStudy (perl):

scroll up to your original integral

OpenStudy (anonymous):

\[next xsex=f(x)(1st function) and (xsinx+cosx)=g(x)(2nd function) and then use ILATE\]

OpenStudy (perl):

stop. look at your original integral above, there is no square term

OpenStudy (anonymous):

yes thats the reason I am not able to solve it

OpenStudy (anonymous):

I solved a different version of the same question...

OpenStudy (perl):

without the square term, it is unsolvable. with the square term, it is solvable

OpenStudy (anonymous):

k......closing the question .......m out!..................THANKS guys!!................Peace..\/

OpenStudy (anonymous):

thats what I figured!!

OpenStudy (perl):

you spent a whole day on an unsolvable integral. here is one integral sqrt(sinx) , good luck

OpenStudy (anonymous):

I know this already

OpenStudy (anonymous):

already wasted my time on this

OpenStudy (anonymous):

the power if in the denominator is 2, then it is easily solvable... And it is not compulsory for every function to have a closed integral .. As easy an integral like integral of sin x/x dx is not possible to have a closed form. @cambrige